DIRC Operations Weekly Report
Period: Wednesday 2007/5/23 - Thursday 2007/5/31
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~8.5 x 1033
(last week: ~8.0 x 1033)
BaBar integrated lumi: 2.81 fb-1
(last weeks: 1.84 fb-1)
Run 6 PEP delivery: 2.88 fb-1
(last weeks: 1.97 fb-1)
Run 6 total delivery: ~37.3 fb-1
Run 6 total BaBar recorded: 35.9 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
5/24: RFS problems. PEP not able to deliver beams for a while.
5/25: Abort gap reduction, three more bunches added to the rings.
5/29: Abort gap still reduced, other two bunches added.
BaBar activities/issues this week:
5/27: Lost 75 min of data taking due to DIRC commissioner error (see DIRC Activities).
- After a global calibration DIRC didn't come to global mode. The pilot was told to go to the DIRC HV control panel
and to put the DIRC in Auto. That fixed the problem.
- Around 3:40pm DIRC Power chiller SIAM red alarm. Physically Chiller and the SIAMs were ok. Ray checked the scope
attached to the chiller power and no glitch was observed.
- Many red alarms about glitches in the power chiller SIAM, some lasted about 4 min. Brad inhibited the SIAM channel via a new Nicolas'
script and added a strip chart in order to monitoring the FEE temperature. Red alarms continued to appear, Brad concluded that Nicolas' script
didn't work, so he inhibited by mistake the wrong SIAM channel manually (the only channel showing a trip, the one that ramp up and down the HV), but no
alarms showed up in the ALH and anything changed in the state machine panel.
- 30min later the DQM noted very distorted DIRC occupancy plots, with very high occupancy in the HV group 11. In consultation with Emmanuele and Nicolas the HV group 1 in sector 0, where there is currently a XTree, was turned off. That didn't fix the problem. We tried to go to Standby to reconfigure but didn't work either. Then Nicolas realized that Brad inhibited by mistake the wrong SIAM channel. The reason of the high occupancy for the HV group in all sectors was that the scalers are in that group and they never ramp down. In consequence we were running at low voltage during stable beams for about 75min (data flagged as unusable).
- The HV SIAM channel was uninhibited and the Pilot and DQM short term instruction updated telling the pilot to check the FEE temperature strip charts and to page the DIRC commissioner in case of any tendency of the temperature going up.
- Nicolas implemented 2 new alarms in the ALH to catch such a scenario if it happens again:
Those alarms give an automatic page to DIRC experts and the guidance tell the pilot how to run scrips for inhibiting the corresponding SIAM channel.
- 1) The SIAM channel status is coherent with the state machine.
- 2) There are more than 5 channels with a not runnable HV.
- Nicolas looked at the DIRC runnable flag and found out that it does not take into account any information about the real hardware
status of the HV channels. That's the reason why nothing happened when the HV SIAM was physically inhibited. Nicolas fixed the problem and
ODC people corrected it according to his directions.
- No more SIAM glitches during the weekend. Power chiller SIAM uninhibited, FEE temperature strip chart removed from pilot console
and Pilot and DQM short term instructions updated.
logbook for details
- Remove scope monitoring chiller if you don't see any glitches.
- Replace two DFBs, one in sector 0 and one in sector 1 with a group of 8 inefficient or flaky channels (require bridge).
- Found a X-tree in sector 7 HV group 55 than only shows up during taking cosmics.
- Found another mini X-tree in sector 0 HV group 1 than only shows up during taking cosmics.
- ROD scheduled for June 21 (~4 hours,extensible to 10), maybe we will replace with Ray the SIAM back card.
Jose, Alejandro and Bradley rotating backup & on-call commissioning duties.